OpenStudy (anonymous):

I nedd to know the correct steps to this question. I know the answer already .

OpenStudy (anonymous):



OpenStudy (anonymous):

What question?

OpenStudy (anonymous):

i jus put it

jhonyy9 (jhonyy9):

write 1=ln10 and if ln(sqrt (x+1))=ln10 than sqrt(x+1)=10

jhonyy9 (jhonyy9):

x+1=100 x=99